OpenStudy (anonymous):

An urn contains three red balls, four white balls, and two black balls. Four balls are drawn from the urn at random without replacement. For each red ball drawn, you win $10, and for each black ball drawn, you lose $15. Let X represent your net winnings. Compute E(X), your expected net winnings.

OpenStudy (anonymous):

probability of red ball=3/9 probability of black ball = 2/9 probability of white ball = 4/9

OpenStudy (dan815):

0

OpenStudy (anonymous):

so how would u compute the expected winnings

OpenStudy (dan815):